OpenStudy (anonymous):

find the slope if it exists of the line containing pair of points (-8,-3) and (-12,-14)

OpenStudy (anonymous):

m = (-14 - 3) / (-12 + 8) = 17/4

OpenStudy (anonymous):

error lol.

OpenStudy (anonymous):

sorry it is a -3, (-14 + 3) / (-12 + 8) = ...

OpenStudy (anonymous):

fruitless so what is the answer

OpenStudy (anonymous):

-11/-4=11/4